Find the standard Gibbs energy charge for reaction



A) CaCO3(s) → CaO(s) + CO2(g)



The ΔGf° values for the three components of this reaction system are CaCO3(s): –1128 kJ mol–1; CaO(s): –603.5 kJ/mol; CO2(g): –137.2 kJ/mol.

1
Expert's answer
2022-03-13T15:17:05-0400

The standard Gibbs energy change for reaction is equal to

"\Delta G_0 (reaction) = \sum n\Delta G_0(products) - \sum m\Delta G_0(reactants)", where n, m -- quantity of moles of each compound.

"\Delta G_0 (reaction) = (-603.5 - 137.2) - (-1128) = 387.3 (kJ/mol)"

Answer: 387.3 kJ/mol is the standard Gibbs energy change of the reaction.